# Encoder Embedding for General Graph and Node Classification

### v1

@@ Abstract
Graph Embedding , General Graph , Asymptotic Theory
t1t1footnotetext: This work was supported in part by the National Science Foundation DMS-2113099, and by funding from Microsoft Research.
[…]
[1]organization=University of Delaware, postcode=19716, city=Newark, country=United States
@@ 1 Introduction
When vertex labels are available, a recent approach known as one-hot graph encoder embedding has demonstrated excellent performance in terms of speed, scalability, and vertex classification [32], making it an ideal choice for handling large graph data compared to other graph embedding techniques. It has also shown effectiveness for multiple-graph and dynamic-graph scenarios [34, 31], and more recently for latent community recovery [33]. However, while there are indications that the method may also be applicable to weighted graphs or distance matrices, its theoretical properties were primarily established for binary graphs.
@@ 2.1 Model Definition
for $j=1,\ldots,m$, and $A_{j}$ must have finite moments. For example, considering an $n\times n$ adjacency matrix $\mathbf{A}$, we can perceive each row of $\mathbf{A}$ (excluding diagonals) as identically (though not necessarily independently) distributed as $\text{Graph}(m,X,\kappa)$.
@@ 2.2 Examples
  • Random Dot Product Graph
    [41, 2]:
@@
  • Stochastic Block Model
    [15, 36]: Let $B=[B(k,l)]\in[0,1]^{K\times K}$ be a block probability matrix. Then
@@
  • Degree-Corrected Stochastic Block Model
    [42]: Let $B=[B(k,l)]\in[0,1]^{K\times K}$ be a block probability matrix, and $\theta_{x}$ and $\theta_{u}$ be two degree parameters for $x$ and $u$ respectively. Then
@@ 2.4 Sample Method
  • Input:
    A general graph matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ and a label vector $\mathbf{Y}\in\{0,1,\ldots,K\}^{n}$, where values from $1$ to $K$ indicate known labels, and the category 00 is used to represent observations with unknown labels.
  • Step 1:
    Calculate the number of known observations per class, denoted as:
@@
  • Step 2:
    Construct the normalized one-hot matrix $\mathbf{W}\in[0,1]^{n\times K}$ as follow: for each observation $i=1,\ldots,n$, set
@@
  • Step 3:
    Perform the following matrix multiplication:
    $\mathbf{Z}=\mathbf{A}\mathbf{W}\in[0,1]^{n\times K}.$
  • Step 4 (Normalization):
    Given $\mathbf{Z}$ from step 3, for each $i$ where $\|\mathbf{Z}(i,\cdot)\|>0$, compute a normalized embedding as follows:
@@
  • Step 5 (Classification Evaluation):
    Let $trn$ be the index set of the training vertices, and $tsn$ be the index set of the testing vertices. Build the linear discriminant analysis model $g(\cdot)$ based on the training data $(\mathbf{Z}(trn,:),\mathbf{Y}_{trn})$. Then compute the classification error by
    $L=(\sum\limits_{i\in tsn}g(\mathbf{Z}(i,:))\neq\mathbf{Y}_{i})/|tsn|$
  • Output:
    The encoder embedding $\mathbf{Z}$ and the cross-validation error $L$.
@@ 3.3 Discriminant Analysis
Given per-class normality, discriminant analysis is the Bayes optimal classifier [9], leading to the following corollary.
Corollary 1.
Suppose Equation 2 holds for every class $y$, then the discriminant analysis is asymptotically Bayes optimal for the classification of $(Z,Y)$, i.e.,
$g(Z)\stackrel{{\scriptstyle n\rightarrow\infty}}{{\rightarrow}}g^{*}(Z)$
where $g(\cdot)$ denotes the discriminant analysis classifier and $g^{*}(\cdot)$ denotes the Bayes optimal classifier.
@@
The Bayes optimal classifier is defined via conditional probability:
$g^{*}(Z)=\arg\max_{y}Prob(Y=y|Z).$
Namely, the theoretical best classifier is based on maximizing the conditional probability. Here,
$Prob(Y=y|Z)=\frac{f_{y}(z)\pi_{y}}{\sum_{k=1}^{K}f_{k}(z)\pi_{k}},$
where $f_{y}(z)$ denotes the class-conditional density of $Z$ given class $y$. Note that the denominator can be omitted since it is constant across classes.
Discriminant analysis exactly equals the Bayes rule when each class-conditional density is normal [9]. Namely, if we assume there exists $h_{y}(\cdot)$, the density for a certain normal distribution, and $f_{y}(z)=h_{y}(z)$, then
$g(Z)=\arg\max_{y}(h_{y}(z)\pi_{y})=g^{*}(Z).$
In our case, we only have asymptotically normal, not exactly normal densities, so the second equality does not hold in the finite sample. However, from Theorem 2, $f_{y}(z)$ is asymptotically normal such that for each $y$,
$f_{y}(z)-h_{y}(z)\rightarrow 0$
for some normal density $h_{y}(z)$. As $\pi_{k}$ is fixed, it follows that
$(f_{y}(z)-h_{y}(x))\pi_{y}\rightarrow 0$
for each $y$, and consequently,
$g^{*}(Z)-g(Z)=\arg\max_{y}f_{y}(x)\pi_{y}-\arg\max_{y}h_{y}(x)\pi_{y}\rightarrow 0$
Therefore, discriminant analysis is asymptotically Bayes optimal for the classification of $(Z,Y)$. In other words, for the encoder embedding, discriminant analysis is asymptotically the Bayes classifier when Equation 2 holds. ∎
@@ 4.1 Simulations
The first scenario involves a high-dimensional Gaussian distribution with the following parameters: $p=100$, $K=3$, and $Y$ taking on the values $\{1,2,3\}$ with equal probability. Within each class $k=1,2,3$, we set $X_{k}|Y=k\sim N(3,1)$, and $X_{d}|Y=k\sim N(1,0.5)$ for $d\neq k$. This setup ensures that the first three dimensions create class separation, while all other dimensions are identical in density across classes. For a given sample size $n$, we generate the sample data $\mathbf{X}$ and corresponding labels $\mathbf{Y}$ based on this multivariate Gaussian setting. We construct the general graph $\mathbf{A}$ by first calculating pairwise Euclidean distances and then applying a distance-to-kernel transformation [27]. Subsequently, we employ the sample encoder embedding, which results in a 3-dimensional representation.
@@ 4.2 Real Data
For each dataset and each method, we conducted 5-fold cross-validation, repeating the process $50$ times. The results, including the average classification error and standard deviation, are summarized in Table 1. Our benchmarks are based on the classification error of the original data whenever available. When the original data is a graph, we use the classification error of the spectral embedding into $30$ dimensions as a benchmark. Finally, we also include the two-layer neural network using MATLAB’s fitcnet method [^1] https://www.mathworks.com/help/stats/fitcnet.html, with a neuron size of $100$ and all parameters set to their default values.
@@ 5 Conclusion
There are many other areas for future works, such as the selection of distance metric in general encoder embedding, incorporation of multiple general graphs, high-dimensional geometric implications, etc., that may lead to further methodological improvements, and enhance our understanding of general graph structure and its relationship to high-dimensional data.

### v2

@@ Abstract
Graph Embedding, General Graph, Asymptotic Theory.
@@ 1 Introduction
When vertex labels are available, a recent approach known as one-hot graph encoder embedding has demonstrated excellent performance in terms of speed, scalability, and vertex classification [31], making it an ideal choice for handling large graph data compared to other graph embedding techniques. It has also shown effectiveness for multiple-graph and dynamic-graph inference [34, 35], graph correlation and hypothesis testing [32], and latent community recovery [33]. However, while there are indications that the method may also be applicable to weighted graphs or distance matrices, its theoretical properties were primarily established for binary graphs.
@@ 2.1 Model Definition
for $j=1,\ldots,m$. For example, considering an $n\times n$ adjacency matrix $\mathbf{A}$, we can perceive each row of $\mathbf{A}$ (excluding diagonals) as identically (though not necessarily independently) distributed as $\text{Graph}(n-1,X,\kappa)$.
@@ 2.2 Examples
  • Random Dot Product Graph [42, 2]:
@@
  • Stochastic Block Model [15, 37]: Let $B=[B(k,l)]\in[0,1]^{K\times K}$ be a block probability matrix. Then
@@
  • Degree-Corrected Stochastic Block Model [43]: Let $B=[B(k,l)]\in[0,1]^{K\times K}$ be a block probability matrix, and $\theta_{x}$ and $\theta_{u}$ be two degree parameters for $x$ and $u$ respectively. Then
@@ 2.4 Sample Method
  • Input: A general graph matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ and a label vector $\mathbf{Y}\in\{0,1,\ldots,K\}^{n}$, where values from $1$ to $K$ indicate known labels, and the category 00 is used to represent observations with unknown labels.
  • Step 1: Calculate the number of known observations per class, denoted as:
@@
  • Step 2: Construct the normalized one-hot matrix $\mathbf{W}\in[0,1]^{n\times K}$ as follow: for each observation $i=1,\ldots,n$, set
@@
  • Step 3: Perform the following matrix multiplication:
    $\mathbf{Z}=\mathbf{A}\mathbf{W}\in\mathbb{R}^{n\times K}.$
  • Step 4 (Normalization): Given $\mathbf{Z}$ from step 3, for each $i$ where $\|\mathbf{Z}(i,\cdot)\|>0$, compute a normalized embedding as follows:
@@
  • Step 5 (Classification Evaluation):Let $trn$ be the index set of the training vertices, and $tsn$ be the index set of the testing vertices. Build the linear discriminant analysis model $g_{trn}(\cdot)$ based on the training data $(\mathbf{Z}(trn,:),\mathbf{Y}_{trn})$. Then compute the empirical classification error by
    $\hat{L}=\frac{\sum_{i\in tsn}1(g_{trn}(\mathbf{Z}(i,:))\neq\mathbf{Y}_{i})}{|tsn|}$
  • Output: The encoder embedding $\mathbf{Z}$ and the classification error $\hat{L}$.
@@ 3.1 Assumptions
The third assumption ensures that the graph weights are well-behaved, a common condition that holds when the raw data have finite moments or when the function $\kappa(\cdot,\cdot)$ is properly normalized.
@@ 3.3 Discriminant Analysis
Given per-class normality, discriminant analysis is the Bayes optimal classifier [9], leading to the following theorem.
Theorem 3.
Suppose Equation 2 holds for every class $y$, and the class-conditional density of the encoder embedding is always bounded. Then, discriminant analysis is asymptotically Bayes optimal for the classification of $(Z,Y)$.
Specifically, given $A\sim\text{Graph}(m,X,\kappa)\in\mathbb{R}^{m}$, and $Z$ as the encoder embedding of $A$, let $g_{m}(\cdot)$ denote the discriminant analysis classifier on $Z$, and $g^{*}_{m}(\cdot)$ denote the Bayes optimal classifier on $Z$. Given any argument $z$ in the support of $Z$, and assuming without loss of generality that $g^{*}_{m}(z)$ always permits a unique solution, it follows that
$g^{*}_{m}(z)-g_{m}(z)\stackrel{{\scriptstyle m\rightarrow\infty}}{{\rightarrow}}0.$
@@
Given any argument $z$ in the support of $Z$, the Bayes optimal classifier is defined as the class that maximizes the conditional probability:
$g^{*}_{m}(z)=\arg\max_{y}Prob(Y=y|Z=z).$
Here,
$Prob(Y=y|Z=z)=\frac{f_{y}^{m}(z)\pi_{y}}{\sum_{k=1}^{K}f_{k}^{m}(z)\pi_{k}},$
where $f_{y}^{m}(z)$ is the class-conditional density of the encoder embedding at graph size $m$. Since the denominator is constant across all possible $y$, it can be safely omitted, leading to:
$g^{*}_{m}(z)=\arg\max_{y}(f_{y}^{m}(z)\pi_{y}).$
Discriminant analysis has the same formulation but, instead of using the true class-conditional density $f_{y}^{m}(z)$, which is usually unknown for sample data, it assumes that all class conditional densities are normal. It then estimates the mean and variance from the sample data. Let $h_{y}^{m}(z)$ denote the normal density used at graph size $m$. The discriminant analysis classifier is given by
$g_{m}(z)=\arg\max_{y}(h_{y}^{m}(z)\pi_{y}).$
Therefore, $g^{*}_{m}(z)$ does not necessarily equals $g_{m}(z)$, because the former uses the true density, while the latter uses an approximated normal density. We aim to prove these two classifiers are asymptotically the same, i.e.,
$g^{*}_{m}(z)-g_{m}(z)\stackrel{{\scriptstyle m\rightarrow\infty}}{{\rightarrow}}0.$
To prove this, consider Theorem 2, which shows that the cumulative distribution of $Z|Y$ is asymptotically normal. Given that the density of the encoder embedding is assumed to be bounded, and since $h_{y}^{m}$ is also bounded as a normal density, with the sample mean and variance estimated under discriminant analysis converging to their true values, we have for each $y$,
$f_{y}^{m}(z)-h_{y}^{m}(z)\rightarrow 0.$
In other words, as the ground-truth density $f_{y}^{m}(z)$ converges to a normal density, it must coincide with the estimated normal density $h_{y}^{m}(z)$ used in discriminant analysis.
Since $\pi_{k}$ is fixed, it follows that
$(f_{y}^{m}(z)-h_{y}^{m}(z))\pi_{y}\rightarrow 0$
for each $y$. Consequently,
$g^{*}_{m}(z)-g_{m}(z)=\arg\max_{y}(f_{y}^{m}(z)\pi_{y})-\arg\max_{y}(h_{y}^{m}(z)\pi_{y})\rightarrow 0.$
It is important to note that while pointwise convergence does not generally imply convergence of the argmax, it does hold in this case. This is because the argmax operation here is applied to a finite and discrete set of $y$ values, and we assume a unique solution for the Bayes optimal classifier with a fixed $z$.
Note that it suffices to assume a unique solution to prove the equivalence. If there are multiple solutions, such that at a given $z$, there are several $y$ that achieves the argmax for $g^{*}_{m}(z)$ as $m$ goes to infinity, it does not matter which $y$ that $g_{m}(z)$ converges to. The discriminant analysis remains asymptotically Bayes optimal by achieving any of the argmax values. Moreover, the classification error remains the same, as long as the maximum conditional density is the same, regardless of how many $y$ values may achieve the argmax. ∎
@@ 4.1 Simulations
The first scenario involves a high-dimensional Gaussian distribution with the following parameters: $p=100$, $K=3$, and $Y$ taking on the values $\{1,2,3\}$ with equal probability. Within each class $k=\{1,2,3\}$, we set $X_{k}|(Y=k)\sim N(3,1)$, and $X_{d}|(Y=k)\sim N(1,0.5)$ for $d\neq k$. This setup ensures that the first three dimensions create class separation, while all other dimensions are identical in density across classes. For a given sample size $n$, we generate the sample data $\mathbf{X}$ and corresponding labels $\mathbf{Y}$ based on this multivariate Gaussian setting. We construct the general graph $\mathbf{A}$ by first calculating pairwise Euclidean distances and then applying a distance-to-kernel transformation [27]. Subsequently, we employ the sample encoder embedding, which results in a 3-dimensional representation.
@@ 4.2 Real Data
For each dataset and each method, we conducted 5-fold cross-validation, repeating the process $50$ times. The results, including the average classification error and standard deviation for the encoder embedding followed by linear discriminant analysis (LDA) or 5-nearest-neighbor (5NN), are summarized in Table 1. We also report three benchmarks: linear discriminant analysis (LDA), 5-nearest-neighbor (5NN), and a two-layer neural network (two-layer NN) using MATLAB’s fitcnet method [^1] https://www.mathworks.com/help/stats/fitcnet.html, with a neuron size of 100 and all parameters set to their default values. These three benchmarks are applied directly to the raw data without the graph transformation.
@@ 5 Conclusion
There are many other areas for future work, such as the selection of distance metrics in general encoder embedding, the incorporation of multiple general graphs, the exploration of high-dimensional geometric implications, and the development of faster computation methods without explicit graph transformation. These directions could lead to further methodological improvements and deepen our understanding of general graph structures and their relationship to high-dimensional data.
Declarations
  • Availability of data and materials: The datasets generated and/or analysed during the current study are available in Github, [https://github.com/cshen6/GraphEmd].
  • Funding: This work was supported in part by the National Science Foundation DMS-2113099, and by funding from Microsoft Research.
  • Acknowledgements: We thank the editor and anonymous reviewers for providing valuable feedback that improved the exposition of the paper.